OpenStudy (anonymous):

Triangle ABC is similar to triangle DEF. If angle A measures 35 degrees and angle E measures 40 degrees, what is the measure of angle C?

OpenStudy (anonymous):

Select one: a. 35 b. 40 c. 75 d. 105

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

180 - 35 - 40 = 105 would be my guess

OpenStudy (anonymous):

i thought u were supposed to add the numbers and see if they aDDED UP TO 180

OpenStudy (anonymous):

well yes but that is what i am doing, say X is the number you want: 180 = 35 (A = D) + 40 (B = E) + X, this gives X = 105 for angles C or F

OpenStudy (anonymous):

ok so is it still 105

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yes, sorry didnt see, i think it is

OpenStudy (anonymous):

ok it better be